OpenStudy (anonymous):

Express in simplest form:x^2-5x/x^2-25 a.x/x-5 b.x/5-x c.-x/x+5 d.x/x+5

OpenStudy (pfenn1):

\[\frac {x^2-5x}{x^2-25}=\frac {x(x-5)}{(x-5)(x+5)}\]Simplify